OpenStudy (anonymous):

WILL GIVE MEDAL!!!!!! A movie club charges a one-time membership fee of $25 which allows members to purchase movies for $7 each. Another club does not charge a membership fee and sells movies for $12 each. How many movies must a member purchase for the cost of the two clubs to be equal?

OpenStudy (deepolisnoob):

You will have 2 equations for this problem For theater one, f(x), you will have a base cost of 25 and a slope of 7, being f(x)=7x+25 For theater two, g(x), you dont have a base, but you have a slope of 12, being g(x)=12x

OpenStudy (deepolisnoob):

you want both theaters to have an equal price, so f(x)=g(x), so 7x+25=12x. You get your x values on one side, making your equation 25=5x. divide both sides by 5, and you are left with x=5

OpenStudy (anonymous):

i think the amount of movies should be 5 because the formula 7x+25=12x. You combine like terms, and the final answer is 5 :)